OpenStudy (anonymous):

Can Someone Help? Part A: Nadia rented a stage platform at $410 for 4 days. If she rents the same stage platform for 8 days, she has to pay a total rent of $770

OpenStudy (anonymous):

Write an equation in the standard form to represent the total rent (y) that Nadia has to pay for renting the stage platform for x days.

OpenStudy (anonymous):

@ganeshie8 You Think you could help?

OpenStudy (the_fizicx99):

Part A: (4,410) and (8,770) 770-410 = 360 = 90 8-4 = 4 y = 90x+50 -90x -90x + y = 50 (-1)-90x + (-1)y = (-1)50 90x - y = -50 <-- standard form of Ax + By = C

OpenStudy (the_fizicx99):

Get it?

OpenStudy (anonymous):

how would i ? Write the equation obtained in Part A using function notation.

OpenStudy (the_fizicx99):

f(x) is the same thing as y, so simply exchange y with f(x). First put it in the form of y = mx + b So, f(x) = 90x + 50

OpenStudy (anonymous):

i get it when people explainn it , but when im on my own, its like im in a space ship

OpenStudy (the_fizicx99):

cx

OpenStudy (anonymous):

Can you help me with the last part?

OpenStudy (the_fizicx99):

Sure!

OpenStudy (anonymous):

Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals

OpenStudy (the_fizicx99):

Hm, well, I would graph the days on the x axis, 0 through 10 with intervals of 1. And graph the cost on the y axis, from 0 through 950 at intervals of 90.